# Numerical linked cluster expansions for two-dimensional spin models with continuous disorder distributions

### v1

@@ V NLCEs in 2D
There are various NLCEs based on different building blocks that have been used in the literature before to study square lattice models. Here we focus on three schemes. First, we introduce a restricted L expansion with a significantly lower number of clusters than the L expansion introduced in Ref. Abdelshafy . We further use the square expansion introduced in Ref. rigol_bryant_07a , and the rectangle expansion introduced in Refs. kallin_13 ; Seing_Rect ; Hazzard_dis . For comparison, we also show some results obtained using the site-based expansion introduced in Ref. rigol_bryant_07a .
The main advantage of the former three techniques is the low number of clusters that enter in the corresponding expansions up to relatively large cluster sizes. This allows us to control the statistical errors of the average over disorder realizations in each cluster, and prevents a significant build-up of those errors that would result in divergences of the bare sums. In the following subsections, we briefly introduce those three NLCE schemes.
@@ V.1 L Expansion
In Ref. Abdelshafy , we developed “strong” and “weak” embedding versions of an NLCE expansion that uses Ls as building blocks. We showed that the strong-embedding version (with all possible Ls connecting the sites present), which requires a smaller number of clusters at each order and hence has a lower computational cost, was preferable as: (i) it has similar convergence properties as the weak embedding version in the high-temperature disordered phase, and (ii) it was the only L expansion that converged when approaching the ground state in ordered phases such as the one in the Ising model. Here, in order to reduce even further the number of clusters of the strong embedding L expansion, we introduce a restricted L expansion. In the restricted expansion, the Ls in the strong embedding NLCE are attached to an existing cluster by sharing the center site, i.e., no L can share only edge sites in any given cluster.
Table 1: Total number of clusters (third column) and the number of topological distinct clusters (second column) in the restricted L and the strong embedding L (in parentheses) expansions vs the number of Ls in the clusters (first column).
No. Ls No. topological clusters Total number of clusters
0 1(1) 1(1)
1 1(1) 1(1)
2 1(2) 2(3)
3 2(6) 5(11)
4 4(18) 13(41)
5 7(61) 34(153)
6 15(202) 90(573)
7 30(700) 239(2162)
8 62(2429) 636(8238)
9 129(8608) 1695(31696)
10 268(30734) 4522(122986)
11 562 12075
12 1178 32265
Such a restriction on the clusters allowed in the strong embedding L expansion results in more compact (larger weight) clusters and reduces the number of clusters significantly. In Fig. 3, we show the clusters with two and three Ls that are present in the restricted L expansion. The total number of clusters [the sum of $L(c)$s in Eq. (3)] at each order of the restricted L expansion is shown in the third column of Table 1. Those numbers are to be compared to the ones in between parentheses, which are the total number of clusters in the strong embedding L expansion Abdelshafy . The second column in Table 1 shows the total number of topologically distinct clusters in each expansion, which are the actual clusters that are diagonalized to compute the observables as they are the ones with different Hamiltonians. One can see that there is an exponential reduction of the number of clusters from the unrestricted to the restricted L expansion as the number of Ls increases. Clusters with the same number of Ls are grouped together and the order $l$ of the expansion is set by the largest number of Ls included in the NLCE sum.
@@ V.2 Square Expansion
Table 2: Sum of $L(c)$s (third column) and the number of topological distinct clusters (second column) in the square expansion vs the number of squares (first column).
No. squares No. topological clusters Sum of $L(c)$s
@@ V.3 Rectangle Expansion
The rectangle expansion was introduced in Ref. kallin_13 to calculate entanglement entropies, and was used in Ref. Seing_Rect to study quench dynamics in clean systems and in Ref. Hazzard_dis to study quench dynamics from inhomogeneous initial states. The rectangle expansion contains clusters that have a rectangular shape. This limits the number of clusters considerably as there are only three possible cluster geometries. For clusters with $N$ sites one can have: (i) a chain of $N$ sites, (ii) a rectangle with $N=N_{x}\times N_{y}$ sites, where $N_{x}$ and $N_{y}$ are the numbers of sites in $x$ and $y$, respectively, for values of $N$ that admit such a decomposition, and (iii) a square with $N=N_{x}^{2}$ sites for $N=4,9,\,\dots$. Squares have $L(c)=1$, while all other clusters have $L(c)=2$, making the combinatorics associated with the rectangle expansion trivial. Clusters with the same number of sites are then grouped together and the order $l$ of the expansion is set by the largest number of sites included in the NLCE sum.
@@ VI Bimodal disorder distribution
In order to gain an understanding of how the L, the square, and the rectangle expansions work in the presence of disorder, in this section we compare our results using those expansions for a bimodal disorder distribution to results obtained in Ref. Tang_BiDisorder using the site expansion. For bimodal disorder, each $J_{\bf ij}$ can have values ±$\pm J$ with equal probability (we set $J=1$). An advantage of such a distribution (e.g., over the continuous ones that we study in the next section) is that one can average over all possible disorder realizations ($2^{b}$, where $b$ is the number of bonds) in the finite clusters considered in the NLCEs so that there are no statistical errors.
@@
Figure 4: Energy per site $E$ vs $T$ for the Ising model with a bimodal disorder distribution. We show results for the restricted L expansion (L) with 7 and 8 Ls, the unrestricted L expansion (L(U)) with 6 and 7 Ls, the site expansion (S) with 13 and 14 sites, and Wynn’s and Euler’s resummations of the site expansion. All the site expansion results are from Ref. Tang_BiDisorder . (Inset) Restricted L expansion with 7 and 8 Ls, and Wynn’s and Euler’s resummations of the site expansion (same results and legends as in the main panel) together with the results for the square expansion ($\square$) with 4 and 5 squares, and the rectangle expansion (R) with 14 and 15 sites.
Recall that the number of clusters in the restricted L expansion grows much more slowly than in the unrestricted one as the number of Ls increases, this is the reason we can compute one order higher of the former expansion for the results shown in Fig. 4. Given the excellent convergence properties of the restricted L expansion, along with the fact that its smaller number of clusters per order will allow to reduce the effect of statistical errors in the subgraph subtractions later when we study continuous disorder distributions, motivate us to focus on that L expansion in what follows. We will refer to the restricted L expansion as the L expansion in the rest of this paper. In the inset in Fig. 4 we compare the Euler and Wynn resummation results for the site expansion with results for the bare sums obtained using the L, the square, and the rectangle expansions. They all agree with each other down to temperatures that are slightly higher than $0.2$, with the L expansion results agreeing with the resummation ones at lower temperatures than the square and rectangle expansions. Having the independent results from the L, the square, and the rectangle expansions will be useful in the rest of this work to gauge convergence for different models.
@@ VII.1 Ising model
Let us first consider the Ising model with the uniform disorder distribution in Eq. (16). For any given cluster $c$, we compute the disorder averages of observables ${\overline{\mathcal{O}(c)}}$ as well as the associated standard deviation $\sigma_{c}({\mathcal{O}})=\sqrt{\overline{\mathcal{O}^{2}(c)}-\overline{\mathcal{O}(c)}^{2}}$. We further compute the normalized standard deviation $\delta_{c}(\mathcal{O})=\sigma_{c}({\mathcal{O}})/{\overline{\mathcal{O}(c)}}$, which is a measure of the statistical errors generated by our finite disorder averages. To show the effect that decreasing $\delta_{c}$ has in our NLCE calculations, in Fig. 6 we plot NLCE results for the energy obtained using the L, square, rectangle, and site expansions (from top to bottom, respectively), when $\delta_{c}(E)\lesssim\varepsilon$ at $T\sim 1$ in all clusters in the expansion. We note that as the size of the clusters increases, because of self-averaging, to achieve the same value of $\delta_{c}(E)$ we need to consider smaller numbers of disorder realizations.
The left (right) panels in Fig. 6 show results when $\varepsilon=5\times 10^{-3}$ ($\varepsilon=10^{-4}$). For $\varepsilon=5\times 10^{-3}$ (left panels in Fig. 6), see Tables 3 and 4 for the number of disorder realizations used, the results for different orders of the expansions differ at temperatures $T>1$, which are sufficiently high for convergence to be achieved given the cluster sizes considered in all the expansions shown, i.e., the lack of convergence observed at those temperatures is a consequence purely of the statistical errors introduced in the averages over disorder realizations. We find the rectangle expansion to be the least affected by the statistical errors for clusters with up to $\sim 15$ sites. This is a result of the simple structure of the subgraph subtraction for this expansion, e.g., as for the 1D expansion considered in Sec. IV.2, there is no accumulation of errors for the chain clusters involved in the rectangle expansion. On the other hand, the site expansion results are strongly affected by the statistical errors even for clusters that have about one-half the number of sites of those in the other expansions.
Table 3: Number of disorder realizations used to obtain the results in Fig. 6 for the L, square ($\square$), and site expansions. The numbers, for $\varepsilon=5\times 10^{-3}$ ($\varepsilon=10^{-4}$, in betwen parenthesis), are shown in units of $10^{3}$ ($10^{6}$).
@@ 12 $10^{4}$ $10^{7}$
For $\varepsilon=10^{-4}$ (right panels in Fig. 6), see Tables 3 and 4 for the number of disorder realizations used, the results for the L, the square, and the rectangle expansions agree at temperatures $T>1$, and are very close to each other at temperatures $0.1\leq T\leq 1$. The site-expansion results, on the other hand, still do not agree with each other at temperatures $T>1$. Since statistical errors of the order of $10^{-4}$ require averages over millions of disorder realizations (see Tables 3 and 4), the lack of convergence of the (low) $7^{\text{th}}$ order of the site expansion makes apparent that such an expansion is not suitable to study models with continuous disorder distributions by averaging over finite numbers of disorder realizations.
In order to improve convergence by reducing the effect of statistical errors even further, we note that the number of subclusters of any given cluster that belongs to order $l$ increases exponentially with $l$, with most of the smallest clusters appearing in the larger clusters. Since the number of clusters also grows exponentially with $l$, the statistical errors of the smallest clusters compound rapidly as the order of the expansion increases. Hence, it is essential to reduce the statistical errors in the smallest clusters as much as possible. To achieve this, in this work, we compute the exact disorder averages for all clusters with up to five sites. Namely, for such clusters, we compute all observables symbolically and then calculate the exact disorder averages by integrating over the continuous disorder distribution. This means that, in what follows, the disorder averages are computed exactly for clusters with one and two Ls in the L expansion, for one square in the square expansion, and for chains with 1 through 5 sites and the square with 4 sites in the rectangle expansion. For higher orders of these expansions, the number of disorder realizations used is about 10 times the ones reported between parentheses in Tables 3 and 4.
Figure 7: Thermodynamic properties of the square lattice Ising model with a uniform disorder distribution. (a) Energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ per site vs $T$ obtained using the L (L), the square ($\square$), and the rectangle (R) expansions. We report results for the highest two orders computed of each NLCE scheme. (Insets) Normalized differences, see Eq. (6), for (a) $E$ at $T=0.01$, (b) $S$ at $T=0.1$, and (c) $C_{v}$ at $T=0.2$ vs $l$ for the same expansions used in the main panels.
@@
The insets in Fig. 7 show how the normalized difference for each observable, see Eq. (6), behaves with increasing the order $l$ of each expansion at a suitably chosen temperature (a temperature at which no differences are visible between the two orders of the NLCE in the main panels). The temperatures chosen are $T=0.01$ [inset in Fig. 7(a)], $T=0.1$ [inset in Fig. 7(b)], and $T=0.2$ [inset in Fig. 7(c)]. In agreement with NLCE results obtained for translationally invariant models in earlier works iyer_15 ; Abdelshafy , one can see that for all observables and all NLCEs considered here the normalized differences decrease exponentially with $l$ before saturating to the statistical errors at large $l$. As expected, given the faster convergence of NLCEs with increasing $T$, we find that the order $l$ at which such a saturation occurs decreases as the temperature increases (not shown).
@@ VII.2 Heisenberg model
Next, we discuss our results for the most challenging model considered in this work. Namely, the square lattice Heisenberg model [see Eq. (2)] with a uniform disorder distribution with zero mean [see Eq. (16)]. This model is frustrated and it is very challenging to study at low-temperature using quantum Monte Carlo simulations because of the sign problem. In Fig. 8, we show results for the highest two orders of the L, the square, and the rectangle expansions for the energy $E$, the entropy $S$, and the specific heat $C_{v}$. For the energy [see Fig. 8(a)], the results for the three expansions are very close to each other down to $T\approx 0.2$, at which $E$ can be seen to begin to plateau to a temperature-independent value. For $S$ [see Fig. 8(b)], all the results are also very close to one another down to $T\approx 0.2$. For $C_{v}$ [see Fig. 8(c)], the results from different expansions depart from each other at temperatures $T\approx 0.4$, below which a maximum appears to develop. Like for the entropy of the Heisenberg model with bimodal disorder in Fig. 5, the L and square expansions are the closest ones for all observables in Fig. 8.
Figure 8: Thermodynamic properties of the square lattice Heisenberg model with a uniform disorder distribution. (a) Energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ per site vs $T$ obtained using the L (L), the square ($\square$), and the rectangle (R) expansions. We report results for the highest two orders computed of each NLCE scheme.
We applied Wynn and Euler resummation techniques to the NLCE results obtained for the square lattice Heisenberg model with a uniform disorder distribution. We found that, for all our observables within the rectangle expansion, Wynn’s algorithm appears to extend the convergence to significantly lower temperatures than the direct sums. Unfortunately, since we have so few orders for the L and the square expansions, none of the resummation algorithms considered extended significantly the convergence of the corresponding direct sums. In Fig. 9, we compare the results of the highest order of the L and the square expansions (6 L and 4 $\square$, respectively) against those obtained for the highest two orders of Wynn’s algorithm applied to the rectangle expansion results.
@@
For the energy [see Fig. 9(a)], the resummation results are very close to each other down to $T=10^{-2}$, and we see a clear plateau for temperatures between $T=10^{-2}$ and $T=10^{-1}$, so we expect the resummation results to be accurate all the way down to the ground state energy. For the entropy [see Fig. 9(b)], the resummation results agree with each other down to $T\approx 0.04$, which is nearly an order of magnitude lower than that at which the direct sums agree with one another. Finally, given the behavior of the direct sums of the rectangle expansion for the specific heat in Fig. 8(c), which appear to develop a maximum at lower temperatures than the L and square expansions, we find the most striking resummation results to be the ones for this observable. The resummation results for the rectangle expansion in Fig. 9(c) depart from those of their corresponding direct sums [shown in Fig. 8(c)] below $T\approx 0.4$, are very close to the direct sums for the square expansion down to $T\approx 0.2$, and very close to each other down to $T\approx 0.06$. The results for the direct sums and the resummations in Fig. 8(c) suggest that a maximum occurs in the specific heat at $T_{m}\approx 0.3$.

### v2

@@ V NLCEs in 2D
There are various NLCEs based on different building blocks that have been used in the literature to study square lattice models. Here we focus on three schemes. First, we introduce a restricted L expansion with a significantly lower number of clusters than the L expansion introduced in Ref. Abdelshafy . We further use the square expansion introduced in Ref. rigol_bryant_07a , and the rectangle expansion introduced in Refs. kallin_13 ; Seing_Rect ; Hazzard_dis . For comparison, we also show some results obtained using the site-based expansion introduced in Ref. rigol_bryant_07a .
The main advantage of the former three NLCEs is the low number of clusters that enter in the corresponding expansions up to relatively large cluster sizes. This allows us to control the statistical errors of the average over disorder realizations in each cluster and prevents a significant build up of those errors that would result in divergences of the bare sums. In the following subsections, we briefly introduce those three NLCE schemes.
@@ V.1 L expansion
In Ref. Abdelshafy , we developed “strong” and “weak” embedding versions of an NLCE expansion that uses Ls as building blocks. We showed that the strong-embedding version (with all possible Ls connecting the sites present), which generally involves a smaller number of clusters at each order and hence has a lower computational cost, was preferable as (i) it has similar convergence properties as the weak embedding version in the high-temperature disordered phase, and (ii) it was the only L expansion that converged when approaching the ground state in ordered phases such as the one in the Ising model. Here, in order to reduce even further the number of clusters of the strong embedding L expansion, we introduce a restricted L expansion. In the restricted expansion, the Ls in the strong embedding NLCE are attached to an existing cluster by sharing the center site, i.e., no L can share only edge sites in any given cluster.
Table 1: Total number of clusters (second and third columns) and the number of topologically distinct clusters (fourth and fifth columns) in the restricted (R) and unrestricted (U) strong embedding L expansions, respectively, versus the number of Ls in the clusters (first column).
No. Ls Total No. clusters No. top. clusters
R U R U
0 1 1 1 1
1 1 1 1 1
2 2 3 1 2
3 5 11 2 6
4 13 41 4 18
5 34 153 7 61
6 90 573 15 202
7 239 2162 30 700
8 636 8238 62 2429
9 1695 31696 129 8608
10 4522 122986 268 30734
11 12075 NA 562 NA
12 32265 NA 1178 NA
Such a restriction on the clusters allowed in the strong embedding L expansion results in more compact (larger weight) clusters and reduces the number of clusters significantly. In Fig. 3, we show the clusters with two and three Ls that are present in the restricted L expansion. The total number of clusters [which for the L expansion equals the sum of $L(c)$s in Eq. (3)] at each order of the restricted L expansion are shown in the second column of Table 1. Those numbers are to be compared to the total number of clusters in the strong embedding L expansion Abdelshafy shown in the third column. The fourth and fifth columns in Table 1 show the total number of topologically distinct clusters in each expansion, which are the actual clusters that are diagonalized to compute the observables as they are the ones with different Hamiltonians. One can see that there is an exponential reduction of the number of clusters from the unrestricted to the restricted L expansion as the number of Ls increases. Clusters with the same number of Ls are grouped together and the order $l$ of the expansion is set by the largest number of Ls included in the NLCE sum.
Among the three main expansions considered in this work, the L expansion is the one that has the most clusters with any given number of sites. This means that the L expansion is the one that best explores the square lattice geometry, and we expect it to provide the most accurate results at intermediate and high temperatures.
@@ V.2 Square expansion
Table 2: Number of topologically distinct clusters (second column) and the sum of $L(c)$s (third column) in the square expansion versus the number of squares (first column).
No. squares No. top. clusters Sum of $L(c)$s
@@ V.3 Rectangle expansion
The rectangle expansion was introduced in Ref. kallin_13 to calculate entanglement entropies, and was used in Ref. Seing_Rect to study quench dynamics in clean systems and in Ref. Hazzard_dis to study quench dynamics from inhomogeneous initial states. The rectangle expansion contains clusters that have a rectangular shape. This limits the number of clusters considerably as there are only three possible cluster geometries. For clusters with $N$ sites one can have (i) a chain of $N$ sites; (ii) a rectangle with $N=N_{x}\times N_{y}$ sites (for values of $N$ that admit such a decomposition), where $N_{x}$ and $N_{y}$ are the numbers of sites in $x$ and $y$, respectively; and (iii) a square with $N=N_{x}^{2}$ sites for $N=4,9,\,\dots$. Squares have $L(c)=1$, while all other clusters have $L(c)=2$, making the combinatorics associated with the rectangle expansion trivial. Clusters with the same number of sites are then grouped together and the order $l$ of the expansion is set by the largest number of sites included in the NLCE sum.
@@ VI Bimodal disorder distribution
In order to gain an understanding of how the L, the square, and the rectangle expansions work in the presence of disorder, in this section we compare our results using those expansions for a bimodal disorder distribution to results obtained in Ref. Tang_BiDisorder using the site expansion. For bimodal disorder, each $J_{\bf ij}$ can have values ±$\pm J$ with equal probability (we set $J=1$). An advantage of such a distribution (e.g., over the continuous ones that we study in the next section) is that one can average over all possible disorder realizations ($2^{b}$, where $b$ is the number of bonds) in the finite clusters considered in the NLCEs so that there are no errors associated to the sampling.
@@
Figure 4: Energy per site $E$ vs $T$ for the Ising model with a bimodal disorder distribution. We show results for the restricted L expansion (L) with 7 and 8 Ls, the unrestricted L expansion [L(U)] with 6 and 7 Ls, the site expansion (S) with 13 and 14 sites, and Wynn’s and Euler’s resummations of the site expansion. All the site expansion results are from Ref. Tang_BiDisorder . Inset: Restricted L expansion with 7 and 8 Ls, and Wynn’s and Euler’s resummations of the site expansion (same results and legends as in the main panel) together with the results for the square expansion ($\square$) with 4 and 5 squares, and the rectangle expansion (R) with 14 and 15 sites.
Recall that the number of clusters in the restricted L expansion grows much more slowly than in the unrestricted one as the number of Ls increases, and this is the reason we can compute one order higher of the former expansion for the results shown in Fig. 4. Given the excellent convergence properties of the restricted L expansion, along with the fact that its smaller number of clusters per order will allow us to reduce the effect of statistical errors in the subgraph subtractions later when we study continuous disorder distributions, we focus on that L expansion in what follows. We will refer to the restricted L expansion as the L expansion in the rest of this paper. In the inset in Fig. 4 we compare the Euler and Wynn resummation results for the site expansion with results for the bare sums obtained using the L, the square, and the rectangle expansions. They all agree with each other down to temperatures that are slightly higher than $0.2$, with the L expansion results agreeing with the resummation ones at lower temperatures than the square and rectangle expansions. Having the independent results from the L, the square, and the rectangle expansions will be useful in the rest of this work to gauge convergence for different models.
@@ VII.1 Ising model
Let us first consider the Ising model with the uniform disorder distribution in Eq. (16). To show the effect that decreasing $\delta_{c}$ [see Eq. (17)] has in our NLCE calculations, in Fig. 6 we plot NLCE results for the energy obtained using the L, square, rectangle, and site expansions (from top to bottom, respectively), when $\delta_{c}(E)\lesssim\varepsilon$ at $T\sim 1$ in all clusters in the expansion. We note that as the size of the clusters increases, because of self-averaging, to achieve the same value of $\delta_{c}(E)$ we need to consider smaller numbers of disorder realizations.
The left (right) panels in Fig. 6 show results when $\varepsilon=5\times 10^{-3}$ ($\varepsilon=10^{-4}$). For $\varepsilon=5\times 10^{-3}$ (left panels in Fig. 6), see Tables 3 and 4 for the number of disorder realizations used, the results for different orders of the expansions differ at temperatures $T>1$, which are sufficiently high for convergence to be achieved given the cluster sizes considered in all the expansions shown, i.e., the lack of convergence observed at those temperatures is a consequence purely of the statistical errors introduced in the averages over finite numbers of disorder realizations. We find the rectangle expansion to be the least affected by those statistical errors for clusters with up to $\sim$15 sites. This is a result of the simple structure of the subgraph subtraction for this expansion, e.g., as for the 1D expansion considered in Sec. IV.2, there is no accumulation of errors for the chain clusters involved in the rectangle expansion. On the other hand, the site expansion results are strongly affected by the statistical errors even for clusters that have about one-half the number of sites of those in the other expansions.
For $\varepsilon=10^{-4}$ (right panels in Fig. 6), see Tables 3 and 4 for the number of disorder realizations used, the results for the L, the square, and the rectangle expansions agree at temperatures $T>1$ and are very close to each other at temperatures $0.1\leq T\leq 1$. The site-expansion results, on the other hand, still do not agree with each other at temperatures $T>1$. Since statistical errors of the order of $10^{-4}$ require averages over millions of disorder realizations (see Tables 3 and 4), the lack of convergence of the (low) 7th order of the site expansion makes apparent that such an expansion is not suitable to study models with continuous disorder distributions by averaging over finite numbers of disorder realizations.
Table 3: Number of disorder realizations used to obtain the results in Fig. 6 for the L, square ($\square$), and site expansions. The numbers, for $\varepsilon=5\times 10^{-3}$ ($\varepsilon=10^{-4}$, in parenthesis), are shown in units of $10^{3}$ ($10^{6}$).
@@ 12 $10^{4}$ $10^{7}$
In order to improve convergence by reducing the effect of statistical errors even further, we note that the number of subclusters of any given cluster that belongs to order $l$ increases exponentially with $l$, with most of the smallest clusters appearing in the larger clusters. Since the number of clusters also grows exponentially with $l$, the statistical errors of the smallest clusters compound rapidly as the order of the expansion increases. Hence, it is essential to reduce the statistical errors in the smallest clusters as much as possible. To achieve this, in this work, we compute the exact disorder averages for all clusters with up to five sites. Namely, for such clusters, we compute all observables symbolically and then calculate the exact disorder averages by integrating over the continuous disorder distribution. This means that, in what follows, the disorder averages are computed exactly for clusters with one and two Ls in the L expansion, for one square in the square expansion, and for chains with one through five sites and the square with four sites in the rectangle expansion. For higher orders of these expansions, the number of disorder realizations used is about 10 times the ones reported between parentheses in Table 3 and in the right column of Table 4.
Figure 7: Thermodynamic properties of the square lattice Ising model with a uniform disorder distribution. (a) Energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ per site vs $T$ obtained using the L (L), the square ($\square$), and the rectangle (R) expansions. We report results for the highest two orders computed of each NLCE scheme. Insets: Normalized differences, see Eq. (6), for (a) $E$ at $T=0.01$, (b) $S$ at $T=0.1$, and (c) $C_{v}$ at $T=0.2$ vs $l$ for the same expansions used in the main panels.
@@
The insets in Fig. 7 show how the normalized difference for each observable, see Eq. (6), behaves with increasing the order $l$ of each expansion at a suitably chosen temperature (a temperature at which no differences are visible between the two orders of the NLCE in the main panels). The temperatures chosen are $T=0.01$ [inset in Fig. 7(a)], $T=0.1$ [inset in Fig. 7(b)], and $T=0.2$ [inset in Fig. 7(c)]. In agreement with NLCE results obtained for translationally invariant models in earlier works iyer_15 ; Abdelshafy , one can see that for all observables and all NLCEs considered here the normalized differences decrease exponentially with $l$ before saturating due to the statistical errors at large $l$. As expected, given the faster convergence of NLCEs with increasing $T$, we find that the order $l$ at which such a saturation occurs decreases as the temperature increases (not shown).
@@ VII.2 Heisenberg model
Next, we discuss our results for the most challenging model considered in this work. Namely, the square lattice Heisenberg model [see Eq. (2)] with a uniform disorder distribution with zero mean [see Eq. (16)]. This model is frustrated and it is very challenging to study at low temperature using quantum Monte Carlo simulations because of the sign problem. In Fig. 8, we show results for the highest two orders of the L, the square, and the rectangle expansions for the energy $E$, the entropy $S$, and the specific heat $C_{v}$. For the energy [see Fig. 8(a)], the results for the three expansions are very close to each other down to $T\approx 0.2$, at which $E$ can be seen to begin to plateau to a temperature-independent value. For $S$ [see Fig. 8(b)], all the results are also very close to one another down to $T\approx 0.2$. For $C_{v}$ [see Fig. 8(c)], the results from different expansions depart from each other at temperatures $T\approx 0.4$, below which a maximum appears to develop. Like for the entropy of the Heisenberg model with bimodal disorder in Fig. 5, the L and square expansions are the closest ones for all observables in Fig. 8.
In the insets in Fig. 8, we contrast the L expansion results for the uniform disorder distribution to those obtained using the same expansion for the clean case. The effect of disorder in the square lattice can be seen to be qualitatively similar to that discussed in chains in the context of Fig. 1. Disorder increases the energies and entropies at all temperatures, as well as displaces the specific heat peak towards lower temperatures and reduces its height. In the presence of disorder, one can also see that the NLCE results for all observables converge at lower temperatures than in the clean case. This highlights a strength of NLCEs for systems with disorder, which in general have shorter correlations. For systems with disorder, NLCEs can provide accurate results to lower temperatures than for their clean counterparts.
Figure 8: Thermodynamic properties of the square lattice Heisenberg model with a uniform disorder distribution. (a) Energy $E$, (b) entropy $S$, and (c) specific heat $C_{v}$ per site vs $T$ obtained using the L (L), the square ($\square$), and the rectangle (R) expansions. We report results for the highest two orders computed of each NLCE scheme. The insets contrast the results for the uniform disorder distribution to the corresponding ones in the clean case. For both sets of results the calculations were done using the L expansion.
We applied Wynn and Euler resummation techniques to the NLCE results obtained for the square lattice Heisenberg model with a uniform disorder distribution. For all our observables within the rectangle expansion, Wynn’s algorithm appears to extend the convergence to significantly lower temperatures than the direct sums. Unfortunately, since we have so few orders for the L and the square expansions, none of the resummation algorithms considered extended significantly the convergence of the corresponding direct sums. In Fig. 9, we compare the results of the highest order of the L and the square expansions (6 L and 4 $\square$, respectively) against those obtained for the highest two orders of Wynn’s algorithm applied to the rectangle expansion results.
@@
For the energy [see Fig. 9(a)], the resummation results are very close to each other down to $T=10^{-2}$, and we see a clear plateau for temperatures between $T=10^{-2}$ and $T=10^{-1}$, so we expect the resummation results to be accurate all the way down to the ground-state energy. For the entropy [see Fig. 9(b)], the resummation results agree with each other down to $T\approx 0.04$, which is nearly an order of magnitude lower than that at which the direct sums agree with one another. Finally, given the behavior of the direct sums of the rectangle expansion for the specific heat in Fig. 8(c), which appear to develop a maximum at lower temperatures than the L and square expansions, we find the most striking resummation results to be the ones for this observable. The resummation results for the rectangle expansion in Fig. 9(c) depart from those of their corresponding direct sums [shown in Fig. 8(c)] below $T\approx 0.4$. Those resummation results are very close to the direct sums for the square expansion down to $T\approx 0.2$, and very close to each other down to $T\approx 0.06$. The results for the direct sums and the resummations in Fig. 8(c) suggest that a maximum occurs in the specific heat at $T_{m}\approx 0.3$.